Operation Summary for May 9, 2011 - May 11, 2011 |
|
Monday May 9 The day shift began with Operations monitoring Tevatron (TEV) store 8719 (61.91E30) and the Recycler stash (358.9E10), with the Antiproton Source (Pbar) stacking (10.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, network storms and problems with a Linac RF station’s (LRF5) low-level RF front end forced a delay in terminating store 8719 until 8:57 AM. Operations put the TEV into HEP shot setup and established store 8720 at 10:35 AM with an average initial luminosity (AIL) of 300.46E30. During the evening shift, Operations terminated beam to T-978 at 6:01 PM. Operations reported some short downtimes due to a MI RF station (MRF13) and due to a quadrupole power supply (E:QD113) breaker trip. Tuesday May 10 The midnight shift began with Operations monitoring store 8720 (82.09E30) and the stash (314.2E10), with Pbar stacking (16.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8720 at 4:03 AM, put the TEV into shot setup, and established store 8721 at 4:13 AM with an AIL of 356.92E30. Operations established beam to T-978 at 6:45 AM. During the day shift, Operations reported that Pbar sweeper magnet controls problems kept stacking off from 1:28 PM to 2:49 PM. During the evening shift, Operations terminated beam to T-978 at 6:04 PM. Operations terminated store 8721 at 8:04 PM. Operations prepared the CDF and D0 collision halls for access; the experiments were out of access by 10:18 PM. Operators turned on the TEV and checked the orbits. Wednesday May 11 The midnight shift began with operators conducting orbit checks, with Operations monitoring the stash (367.0E10), with Pbar stacking (7.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TEV into shot setup at 1:07 AM and established store 8725 at 2:20 AM with an AIL of 332.28E30. Operations attempted to established beam to T-978, but at 4:20 AM a Switchyard power supply (S:HP3US) failed at 4:20 AM. Operators called in EE Support personnel who repaired the supply. Operations established beam to T-978 at 7:16 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow D0 a collision hall access after store 8725 is terminated later today. She will also allow MI personnel to conduct a coalescing check and Booster personnel to conduct a study. |
|
More Information Comments and Suggestions |
| last modified 05/11/2011 email Fermilab |
| Security, Privacy, Legal | |